OpenStudy (mackenzie2013):

HELP! (11-2)! 4!- ------- 8!

OpenStudy (mackenzie2013):

Can someone explain with steps?!

OpenStudy (michele_laino):

we have the subsequent steps: \[\large \left( {4!} \right) \times \frac{{\left( {11 - 2} \right)!}}{{8!}} = 24 \times \frac{{9!}}{{8!}} = 24 \times \frac{{\left( {8!} \right) \times 9}}{{\left( {8!} \right)}} = 24 \times 9 = ...\]

OpenStudy (michele_laino):

since: n! = (n-1)!*n

OpenStudy (michele_laino):

and: 4!=4*3*2*1= 24

OpenStudy (mackenzie2013):

Thank you!